# Comparing Bad Apples to Good Oranges: Aligning Large Language Models via Joint Preference Optimization

### v1

@@ Abstract
A common technique for aligning large language models (LLMs) relies on acquiring human preferences by comparing multiple generations conditioned on a fixed context. This only leverages the pairwise comparisons when the generations are placed in an identical context. However, such conditional rankings often fail to capture the complex and multidimensional aspects of human preferences. In this work, we revisit the traditional paradigm of preference acquisition and propose a new axis that is based on eliciting preferences jointly over the instruction-response pairs. While prior preference optimizations are designed for conditional ranking protocols (e.g., DPO), our proposed preference acquisition protocol introduces Dove, a new preference optimization objective that upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair. Interestingly, we find that the LLM trained with joint instruction-response preference data using Dove outperforms the LLM trained with DPO by $5.2\%$ and $3.3\%$ win-rate for the summarization and open-ended dialogue datasets, respectively. Our findings reveal that joint preferences over instruction and response pairs can significantly enhance the alignment of LLMs by tapping into a broader spectrum of human preference elicitation. The data and code is available at https://github.com/Hritikbansal/dove.
@@ 1 Introduction
Recently, alignment [42, 31] has emerged as a crucial step in enhancing the performance of large language models (LLMs) [1, 30, 44, 3, 11, 48, 19] in diverse real-world applications [22, 65, 57, 12, 21]. In particular, the aligned LLMs can produce high-quality outputs that satisfy human values and intents, including helpfulness, coherence, and harmlessness [4, 31]. LLM alignment hinges on the quality of preferences acquired from human or AI annotators [31]. Among the various preference acquisition protocols [23, 58, 39, 8], the ranking-based approach is the most widely used paradigm for aligning LLMs[42, 31, 6, 49, 45]. Specifically, in this approach the annotator has to compare a pair of responses conditioned on a fixed context. For instance, human can select a ‘preferred’ response by comparing a pair of responses for the instruction ‘Create a list of four fruits other than Apple’ (Figure 1 (left)).
Besides ranking preferences conditioned on a fixed context, humans can also express preferences in non-identical contexts. For example, while browsing reviews for products on an e-commerce website, humans are likely to prefer an accurate and detail-oriented review for a camera over an incoherent, vague movie review even though the products (camera and movie) are qualitatively different. Although the traditional conditional rankings provide rich preference for alignment, they fail to holistically capture the various dimensions of reasoning of human preferences. In this work, we revisit the traditional paradigm of conditional preference acquisition and propose a new approach for jointly eliciting preferences over instruction-response pairs. This method aims to uncover diverse reasoning paths in the process of acquiring feedback.
In this work, we develop a framework to acquire preferences jointly over instruction-response pairs. Starting from an instruction-response data consisting of response $R_{i}$ for instruction $I_{i}$ (say $i\in\{1,2\}$), we acquire ranking-based preferences over the instruction-response pairs $(I_{1},R_{1})$ and $(I_{2},R_{2})$. As shown in Figure 1 (right), we aim to understand whether the response in the pair $X$ is perceived better than the response in the pair $Y$. For instance, humans would prefer a helpful response to the instruction ‘Create a list of four fruits’ over a response that completely ignores the instruction ‘Create a list of beach activities’. This suggests that we can reveal preference axes like adherence to instructions, grammatical fluency, and clarity, even when following a joint preference optimization protocol. In addition, our protocol can elicit human preference behaviours that are obfuscated in the prior protocols, and redefines conditional preference elicitation as a special case where the instructions are identical.
@@
Figure 1: Overview of the proposed Joint Preference Optimization paradigm. (Left) We show that the conditional preference acquisition method would require the annotators to compare two responses for an identical instruction. (Right) We show that the annotators can also assign rankings jointly over instruction-response pairs. Specifically, the annotator prefers a helpful response (e.g., Apple … Grape) over a response that ignores the context of the instruction (e.g., wear sunscreen … litter). Our framework thus elicits preferences that are obfuscated in the prior approach.
Next, we propose Dove, a framework for aligning LLMs with our proposed joint preference elicitation scheme. Specifically, it upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair. This differs from the other frameworks that assume conditional rankings in their feedback data, such as DPO [35, 5] and preference optimizations that train a separate reward model such as PPO and rejection sampling [40, 29]. Further, we further point that Dove subsumes the prior preference optimizations as conditional rankings are a special case of joint preferences (e.g., when $I_{1}=I_{2}$). In our experiments, we focus on extending and comparing against DPO because of their simplicity, stability, and high-performance. However, our framework can be easily applied to PPO-like approaches by training a reward model on the joint preferences.
@@
Finally, we conduct experiments to explore the new reasoning paths enabled by joint preference elicitation, followed by aligning LLMs with the Dove objective. To do so, we explore the interplay between the feedback data collected under conditional rankings and joint preferences protocol. In addition, we ask human annotators to explain their preference decisions, uncovering new reasoning paths that highlight the complexities of the preference acquisition process (§4). After feedback acquisition, we aim to investigate the impact of diverse preferences collected from conditional and joint preferences on LLM alignment. In our experiments, we align a Mistral-7B LLM with the preferences acquired from the conditional rankings and joint preferences, using our Dove algorithm. We find that the Dove outperforms the supervised finetuned LLM by $30\%$ and $18\%$ win-rate against the gold responses on the unseen instructions from the summarization and open-ended dialogues datasets, respectively. Surprisingly, we find that Dove can effectively tap into the diverse preferences in the conditional and joint feedback data and outperforms DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and open-ended dialogues, respectively. This indicates that by utilizing the diverse preference signals present in the existing data, we can align an LLM robustly without acquiring additional instruction-response data.

### v2

@@ Abstract
A common technique for aligning large language models (LLMs) relies on acquiring human preferences by comparing multiple generations conditioned on a fixed context. This method, however, relies solely on pairwise comparisons, where the generations are evaluated within an identical context. While effective to such conditional preferences often fail to encompass the nuanced and multidimensional nature of human preferences. In this work, we revisit the traditional paradigm of preference acquisition and propose a new axis based on eliciting preferences jointly over the instruction-response pairs. Unlike prior preference optimizations, which are designed for conditional ranking protocols (e.g., DPO), we propose Joint Preference Optimization (JPO), a new preference optimization objective that upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair. Interestingly, LLMs trained with joint instruction-response preference data using JPO outperform LLM trained with DPO by $5.2\%$ and $3.3\%$ win-rate for summarization and open-ended dialogue datasets, respectively. Our findings reveal that joint preferences over instruction and response pairs can significantly enhance the alignment of LLMs by tapping into a broader spectrum of human preference elicitation. The data and code is available at https://github.com/Hritikbansal/dove.
Comparing Bad Apples to Good Oranges:
Aligning Large Language Models via Joint Preference Optimization
Hritik Bansal♡∗      Ashima Suvarna♡∗      Gantavya Bhatt♢∗ Nanyun Peng     Kai-Wei Chang      Aditya Grover University of California, Los Angeles University of Washington, Seattle {hbansal,asuvarna31}@cs.ucla.edu,{gbhatt2}@uw.edu
@@ 1 Introduction
Recently, alignment (Stiennon et al., 2020; Ouyang et al., 2022) has emerged as a crucial step in enhancing the performance of large language models (LLMs) (Anthropic, 2024; OpenAI, 2023; Team et al., 2023; Anthrophic, 2023; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023) in diverse real-world applications (Li et al., 2023; Zheng et al., 2023a; Wu et al., 2023a; Clusmann et al., 2023; Lambert et al., 2024). In particular, aligned LLMs generate responses that maximize human utility along various dimensions such as helpfulness, coherence, and harmlessness (Askell et al., 2021; Ouyang et al., 2022). Here, the notion of human utility is subjective Kirk et al. (2024); Gabriel (2020), and mainly hinges on how preferences are acquired from annotators Otto et al. (2022). Among the various preference acquisition protocols (Lightman et al., 2023; Wu et al., 2023b; Scheurer et al., 2023; Bansal et al., 2023), the ranking-based approach is the most widely used paradigm to align LLMs (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022a; Tunstall et al., 2023; Teknium, 2023). Specifically, in ranking approach the annotator has to compare a pair of responses conditioned on a fixed context. For instance, humans can select a ‘preferred’ response by comparing a pair of responses for the instruction ‘Create a list of four fruits other than Apple’ (Figure 1 (left)).
@@
Figure 1: Overview of the Joint Preference Optimization. (Left) We show that the conditional preference acquisition method would require the annotators to compare two responses for an identical instruction. (Right) We show that the annotators can also assign rankings jointly over instruction-response pairs. Specifically, the annotator prefers a helpful response (e.g., Apple … Grape) over a response that ignores the context of the instruction (e.g., wear sunscreen … litter). Our framework thus elicits preferences that are obfuscated in the prior approach.
Although traditional conditional rankings provide rich preferences for alignment, they fail to holistically capture the multi-faceted nature of human decision-making and preferences (Zhi-Xuan et al., 2024; Gigerenzer, 2008). Besides ranking preferences conditioned on a fixed context, humans can also express preferences in non-identical contexts. For example, while browsing reviews for products on an e-commerce website, humans are likely to prefer an accurate and detail-oriented review for a camera over an incoherent, vague movie review even though the products (camera and movie) are qualitatively different.
In this work, we revisit the traditional paradigm of conditional preference acquisition by developing a framework to acquire preferences jointly over instruction-response pairs. Starting from an instruction-response data consisting of response $R_{i}$ for instruction $I_{i}$ (say $i\in\{1,2\}$), we acquire ranking-based preferences over the instruction-response pairs $(I_{1},R_{1})$ and $(I_{2},R_{2})$. As shown in Figure 1 (right), we aim to understand whether the response in the pair $X$ is perceived better than the response in the pair $Y$. We hypothesize that by capturing preferences in non-identical contexts our protocol can elicit human behaviors that are obfuscated in prior protocols. First, wee show that humans can provide decisive preferences in joint preferences protocol (5.4). Then, we analyze how joint preferences differ from conditional preferences on the same dataset(§5.4).
Preference Acquisition Algorithm Alignment Objective Different Instructions
Score Ethayarajh et al. (2024) Conditional No
Comparison (DPO Variants) Rafailov et al. (2024) Conditional No
Park et al. (2024) Conditional No
Liu et al. (2024) Conditional No
Meng et al. (2024) Conditional No
Hong et al. (2024) Conditional No
Pairwise JPO (ours) Joint Yes
Table 1: JPO differs from prior works along three key aspects: preference acquisition (scoring or comparison), objective (conditional or joint distribution), and their ability to handle non-identical instruction-responses.
Prior works like DPO and its variants Rafailov et al. (2023); Yin et al. (2024); Liu et al. (2024); Meng et al. (2024); Hong et al. (2024); Azar et al. (2023) rely on conditional rankings, and thus do not have access to the joint distribution of human preferences in the ranking protocol (Table 1). While a rating protocol Ethayarajh et al. (2024) allows for a comparison between responses from non-identical instructions, it can be inconsistent with rankings Bansal et al. (2023) and ignores the possibility of preferences over a pair of chosen or rejected responses.  [^1] For instance, a pair of responses that achieves a score of 0, under the rating protocol, will result in an indecisive preference. Thus, we propose Joint Preference Optimization (JPO), a framework for aligning LLMs with our proposed joint preference elicitation scheme. Specifically, it upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair. Furthermore, JPO subsumes prior preference optimizations as conditional rankings are a special case of joint preferences (e.g., when the instructions are identical).
@@
We conduct experiments to explore new reasoning paths enabled by joint preference elicitation and alignment of LLMs with the JPO objective. By analyzing feedback from conditional rankings and joint preferences protocols, along with explanations from human annotators, we uncover the complexities of the preference acquisition process (§5). Using our JPO algorithm, we align a Mistral-7B LLM with the collected preferences, achieving a $30\%$ and $18\%$ higher win rate against gold responses on unseen instructions from summarization and dialogue datasets, respectively. JPO leverages diverse preferences effectively, outperforming DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and open-ended dialogues, respectively. In addition, JPO outperforms KTO by $3.5\%$ on the open-ended dialogues dataset. It also surpasses both DPO and KTO in the AlpacaEval2 benchmark (§6). This indicates that by utilizing the diverse preference signals present in the existing data, we can align an LLM robustly without acquiring additional instruction-response data.
2 Related Work
Alignment using Reinforcement Learning.
Aligning LLMs with human preferences using reinforcement learning is widely adopted to ensure LLMs follow user intents without being harmful Ouyang et al. (2022). This alignment is usually done by first optimizing for a reward model on preference data Bradley and Terry (1952); Likert (1932); Bansal et al. (2023), followed by aligning the LLMs distribution that maximizes the learned reward model using Reinforcement Learning (RLHF) Schulman et al. (2017); Ouyang et al. (2022), with optional Divergence penalty Wang et al. (2023a) to avoid deviating from the reference policy. Additionally, Dubois et al. (2023); Lu et al. (2024); Zheng et al. (2023b) observe that preferences from LLMs can also be used for alignments motivating Reinforcement Learning through AI feedback (RLAIF). Contrary to prior work that collect preferences as conditional rankings, we emphasize that preference acquisition is a complex phenomenon and elicit joint preferences over instruction-response data.
Reward Free Policy Alignment.
Rafailov et al. (2024) introduced Direct Preference Optimization (DPO) that optimizes directly within the model parameter space, hence eliminating the reward modeling step. Liu et al. (2024) extends this framework where instead of two responses, alignment is done over the list of responses while Liu et al. (2023) improves DPO using statistical rejection sampling. Amini et al. (2024) provides an offset in the DPO objective to increase the margins and Pal et al. (2024) suggests adding an explicit penalty term to avoid a reduction in the likelihood of preferred pairs over the DPO training. Recent variants of DPO such as SimPO (Meng et al., 2024) alleviates the need of reference policy in the objective. Contrary to our work where we compare the joint distributions, Yin et al. (2024) proposes RPO that compares the conditional likelihood of a winning response with the losing response of another prompt. Beyond DPO, Ethayarajh et al. (2024) proposed a human-aware loss function-based framework using prospect theory named KTO, and Azar et al. (2023) proposes IPO that uses human preferences expressed as pairwise preferences. Lastly, Zhao et al. (2022) uses sequence likelihood calibration to align the model from human preference. Despite of a vast body of work arising from DPO, none of the existing methods can operate and contrast over the joint distribution of instruction-response pairs like the proposed JPO algorithm.